OpenStudy (anonymous):

Jo Ann white is paid a salary of 410$ a week and a commission of 5.6% on all sales . Her sales last week were 6700$ . Find her total earnings for the week.

OpenStudy (campbell_st):

Its simply \[$410 + 6700\times \frac{5.6}{100}\]

OpenStudy (perl):

Jo Ann white is paid a salary of 410$ a week and a commission of 5.6% on all sales . Her sales last week were 6700$ . Find her total earnings for the week. J = 410 + .056*x, where x is total sales. J = 410 + .056 * 6700